Circular No. 8783

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 T10 AND C/2006 U8–U11 (SOHO) Additional apparent comets have been found on SOHO website images

(cf. IAUC 8780), all being Kreutz sungrazers except C/2006 U10 (Meyer group). K. Battams notes that C/2006 T10 was small and extremely faint (mag ∼ 8.5). C/2006 U8 was tiny, stellar, and faint (mag ∼ 8.0) in C3 images, and extremely faint and diffuse in C2 images. C/2006 U9 was small and stellar in appearance (mag ∼ 7) with big data gaps in C3 images, but slightly diffuse (mag 6.5–7) with a hint of a short tail in C2 images. C/2006 U10 was stellar in appearance and very faint, fading out as it neared the sun. C/2001 U11 appeared stellar and reached mag ∼ 6.5 in C3 images, while it showed a stubby tail in C2 images.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 T10 Oct. 12.538 13^{h}03.^{m}0 − 8^{o}15′ C2  BZ 2006-X25
C/2006 U8  16.532 13 14.8 −11 01 C3/2 TS 2006-X25
C/2006 U9  19.679 13 20.4 −11 47 C3/2 SF 2006-X25
C/2006 U10 20.104 13 34.6 − 8 45 C2 TH 2006-X26
C/2006 U11  24.138 13 36.1 −13 37 C3/2 HS 2006-X26

COMET C/2006 X1 (LINEAR) An apparently asteroidal object discovered by the LINEAR survey, and

posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been found to have a coma of diameter 12′′ surrounding a central condensation of diameter 5′′ on CCD images taken on Dec. 11.94–11.99 UT by P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m Schmidt-Cassegrain reflector). A. R. Gibbs reports that four co-added 120-s unfiltered CCD images taken in poor seeing with the Mt. Lemmon 1.5-m reflector on Dec. 12.2 show a compact coma and a narrow 90′′ tail in p.a. 260^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Dec. 11.12465  0^{h}10^{m}06.^{s}42  +41^{o}10′59′′.8  19.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-X41.

T = 2007 June 24.183 TT  ω = 141.^{o}659
Ω = 259.330 }2000.0
q = 6.01297 AU  i = 40.731
2006 December 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green